Grade 8General Physics

a wheel rotating with uniform angular acceleration covers 50 revolutions in the first five second after the start . find the angular acceleration and the angular velocity at the end of five seconds.

prasanjeet kumar
12 Years agoGrade 8

2 Answers

neha
9 Years ago
Is q m sbse pehle hum angular acceleration nikalenge wo 4 aayega equation use krk usk baad isme hum angular velocity nikalenge equation se w=w•+alpha×t se iska answer 40Π aajayega
Prashant
8 Years ago
Α = angular acceleration = rad/sec²t = time = 5 sec.ω₀ = initial angular velocity = 0 rad/sec as the body is at rest at t = 0 sec.θ = total angle rotated from t=0 to t, = 50 * 2π radians.Total angle rotated = θ = ω₀ t + 1/2 * α * t² (It`s like: total distance = s = u t + 1/2 * a * t²)So 100 π = 0 * 5 + 1/2 * α * 5² α = 8π rad/sec² = 8π /(2π) revolutions/sec² , as 1 revolution = 2π rad. = 4 revolutions/sec². α = 4 rev/sec.========================= direct calculation in rev/sec²Total angle rotated = θ = ω₀ t + 1/2 * α * t² 50 rev = 0 * 5 + 1/2 * α * 5² α = 4 rev/sec²